# Dynamical phase transitions in the non-reciprocal Ising model

### v1

@@ V.1 Connection with locally-coupled oscillators
Equations (25)-(27) can be viewed as describing locally-coupled oscillators with internal frequency and discrete states. In this sense, they share similarities with the Wood et al. cyclic model [91, 90, 92] albeit in that model oscillators are restricted to forward moves (backward moves are prohibited). For $K=0$, our model reaches an equilibrium state that is equivalent to the clock model (vector Potts model) for $q=4$ [148, 149].
Fig. 3: Mean-field (MF) flow in $\left(m_{A},m_{B}\right)$ space. Arrows correspond to the vector $\vec{v}=(\partial_{t}m_{A},\partial_{t}m_{B})$ calculated from the expanded MF equation (top row, Eq. (30)) and full MF equation (bottom row, Eq. (24) with $D=0$). The limit cycle in the swap phase is highlighted by a semi-transparent blue line. Non-reciprocal inter-species coupling is $\tilde{K}=0.4$, and reciprocal coupling between same-species nearest neighbours is (a,d) $\tilde{J}=0.8$ , (b,e), $\tilde{J}=1.3$, and (c,f) $\tilde{J}=1.8$, showing the disorder, swap, and static-order phase, respectively. While the exact flow varies between expanded (top) and full (bottom) MF equations, the qualitative behavior is the same.
@@ V.2 Analysis of the bifurcations
In the ${m_{\alpha}\ll 1}$ limit and for small gradients, the rescaled MF equation in the fully anti-symmetric case (Eq. (24)) can be approximated by
@@
For $\tilde{K}=0$, Eq. (28) reduces to two uncoupled MF equations associated with $\phi^{4}$ Ginzburg-Landau theory. For any $\tilde{K}\neq 0$ the equations can not be derived from a Hamiltonian dynamics nor a gradient descent, which corresponds to its far-from-equilibrium nature. Note that due to the lack of rotational symmetry in the cubic term, Eq. (28) also differs from the Complex Ginzburg-Landau (CGL) equation [84, 150] given by
@@
where $\psi({\vec{r}},t)$ is a complex field and $a$, $b$ and $c$ are complex numbers. However, Eq. (28) becomes CGL-like in special limits as discussed in Sec. V.2.
A linear stability analysis shows that any homogeneous solution, $m_{\alpha}({\vec{r}},t)\equiv m_{\alpha}(t)$, that constitutes a stable fixed point of the spatially homogeneous case ($D=0$) remains stable in the inhomogeneous case ($D>0$). Therefore, in what follows we analyze the spatially homogeneous MF equation given by
@@
$-\frac{1}{3}\left(\tilde{J}m_{\alpha}+\tilde{K}\varepsilon_{ab}m_{\beta}\right)^{3}.$
Figure 3 shows the phase-space velocity field $\vec{v}=(\partial_{t}m_{A},\partial_{t}m_{B})$ from Eq. (30) along with the fixed points which are obtained by numerically solving the $\partial_{t}m_{A}=\partial_{t}m_{B}=0$ equations. The stability of the fixed points (shown in the figure as orange, pink, and brown for stable, saddle, and unstable, respectively) is determined from the eigenvalues of the Jacobian matrix $\mathcal{J}$ of the linearized equation $\partial_{t}\delta\vec{m}=\mathcal{J}\delta\vec{m}$ where linearization is performed around the fixed point.
@@
The MF phase diagram is shown in Fig 4, as a function of the couplings $\tilde{J}$ and $\tilde{K}$. The phases are determined by the fixed points and their stability as described above, but for generality sake we used the full non-linear homogeneous equation, namely Eq. (24) with $D=0$. When non-reciprocal interactions are turned off (${\tilde{K}=0}$), a pitchfork bifurcation at ${\tilde{J}_{c}\equiv 1}$ (red point in Fig. 4) separates the disordered phase (in blue) from the static-order phase (in green), as in the usual Ising model. When non-reciprocal interactions are present (${\tilde{K}\neq 0}$), the swap phase arises (in pink) and is separated from the disordered state by a Hopf bifurcation at $\tilde{J_{c}}=1$ (green line), and from the static-order phase by another bifurcation known as a saddle-node on an invariant circle (SNIC) bifurcation [82, 64] (blue line ${\tilde{K}_{\rm SNIC}(\tilde{J})}$). We analyze each one of these bifurcations below.
Fig. 4: Mean-field phase diagram. The three phases shown on the diagram are (i) disorder, (ii) swap and (iii) static order. The lines separating the phases are Hopf bifurcation (thin green) and SNIC bifurcation (thick blue). Red dot is the Pitchfork bifurcation. Obtained from numerical analysis of Eq. (24) with $D=0$.
@@ V.2.1 Pitchfork bifurcation
For ${\tilde{K}=0}$, Eq. (30) leads to the standard pitchfork bifurcation at ${\tilde{J}_{\rm PF}=1}$. The ${m_{A}=m_{B}=0}$ point becomes unstable and four stable fixed points
$m_{A}=\pm m_{B}=\pm[3(\tilde{J}-1)/\tilde{J}^{3}]^{1/2}$ (31)
@@ V.2.2 Hopf bifurcation
This means that there is a Hopf bifurcation at ${\tilde{J}_{\rm Hopf}=1}$: the ${m_{A}=m_{B}=0}$ point become unstable and a stable limit cycle emerges, corresponding to the swap phase.
The limit cycle is better understood in polar coordinates, ${m_{A}=r\cos\theta}$ and ${m_{B}=r\sin\theta}$ where Eq. (30) becomes
$\partial_{t}r$ $=-\left(1-\tilde{J}\right)r-\frac{1}{4}\tilde{J}^{3}\bigg{[}1+\kappa^{2}$
@@
$+\left(\frac{\kappa^{3}}{3}-\kappa\right)\cos(4\theta)+\left(\frac{1}{3}-\kappa^{2}\right)\sin(4\theta)\bigg{]}r^{2}\;,$ (34)
@@
The limit cycle leads to anharmonic oscillations, where both $r$ and $\dot{\theta}$ depend on $\theta$. Its shape has a rotational symmetry of order 4 due to the $C_{4}$ symmetry of the mean-field equation. However, very close to the bifurcation, i.e. $\tilde{J}-1\ll 1$, the oscillations are harmonic [151]. This can be seen as follows.
@@
where we used the fact that ${r^{2}\sim 1-\tilde{J}}$. The angle $\theta$ then has the solution ${\theta=\theta_{0}+\tilde{K}t}$, and the equation for $r$ becomes
$\partial_{t}r$ $=-\left(1-\tilde{J}\right)r-\frac{1}{4}\bigg{[}1+\tilde{K}^{2}+\left(\frac{1}{3}-\tilde{K}^{2}\right)\cos(4\tilde{K}t)$ (36)
@@
$r\left(t\right)=2\sqrt{\frac{\tilde{J}-1}{1+\tilde{K}^{2}}}\times\frac{1}{\sqrt{1-\left(\tilde{J}-1\right)f\left(\tilde{J},\tilde{K},t\right)}}$ (37)
@@
$f\left(\tilde{J},\tilde{K},t\right)$ $=\frac{1-\tilde{J}+3\left(\tilde{J}-3\right)\tilde{K}^{2}+2\tilde{K}^{4}}{3\left(1+\tilde{K}^{2}\right)\left[\left(\tilde{J}-1\right)^{2}+4\tilde{K}^{2}\right]}\cos\left(4\tilde{K}t\right)$ (38)
$-\frac{\tilde{K}\left(5-3\tilde{J}+\left(\tilde{J}-7\right)\tilde{K}^{2}\right)}{3\left(1+\tilde{K}^{2}\right)\left[\left(\tilde{J}-1\right)^{2}+4\tilde{K}^{2}\right]}\sin\left(4\tilde{K}t\right)$
@@
which is a circular motion of constant angular velocity in phase space. This harmonic solution can be rationalized by looking at the two times scales corresponding respectively to the angular and radial motion close to the bifurcation (described by Eqs. (35) and (36), respectively). The time scale for the angular motion is $T_{\theta}\sim\tilde{K}^{-1}$ while the time scale of the radial motion is determined by the relaxation time of $r(t)$ to its new minimum when $\theta$ changes. Thus, it is found by linearizing the equation for $r$ around 00: this yields $T_{r}\sim(\tilde{J}-1)^{-1}$. The two time scales $T_{\theta}$ and $T_{r}$ can be viewed as “oscillation time-scale” and “alignment time-scale”, respectively. Since $T_{r}\gg T_{\theta}$, the angular motion is much faster and there is not enough time for $r$ to reach a new minimum when $\theta$ changes. Thus, only the $\theta$-averaged value of the right hand side of Eq. (V.2.2) matters. This is supported by Fig. 5 which shows typical phase-space trajectories in the swap phase: as $\tilde{J}-1$ increases, the motion indeed loses its rotational symmetry and become more “square-like”.
Fig. 5: Swap phase at different distances from the Hopf bifurcation. Numerical solutions of the spatially homogeneous MF dynamics, Eq. (30). (a-c) $m_{B}$ vs. $m_{A}$. (d-f) $m_{A}$ vs. $t$. From left to right: $\tilde{J}=1.0002$, $1.001$, and $1.0028$. In all plots $\tilde{K}=0.001$. Red dot is the initial state and dashed black line is the steady state predicted from Eq. (39), valid in the $\tilde{J}-1\ll 1$ limit.
The above result is a consequence of a more general theorem. A reaction-diffusion equation that undergoes a Hopf bifurcation such as the MF evolution Eq. (24) can be reduced, close to the bifurcation, to the CGL equation (i.e., the CGL equation is its normal form). The transformation to the CGL equation guarantees that the steady-state solution of the non-transformed set of equations is an elliptical limit cycle (or in special cases like here, circular) [151]. Because of this reduction, we expect the disorder-to-swap phase transition in our model to exhibit the same critical exponents as the Hopf bifurcation in the CGL equation, which has been studied in Refs. [85, 86].
@@ V.2.3 Infinite-period (SNIC) bifurcation
Let us now start from the limit-cycle swap phase. Upon increasing $\tilde{J}$ or decreasing $\tilde{K}$, the system goes through a ($C_{4}$ equivariant) saddle-node on an invariant circle (SNIC) bifurcation, where four pairs of stable and saddle fixed points emerge from the limit cycle (pink and orange points in Fig. 3c,e). This bifurcation is also known under several other names: infinite-period bifurcation (because the period of oscillation diverges at the bifurcation point), saddle-node on a limit cycle (SNLC), saddle-node infinite-period (SNIPER), etc. Once the stable fixed points emerge, the system is in a static-order phase.
Noticing that the infinite-period bifurcation occurs when $\tilde{K}$ and $\tilde{J}-1$ are comparable (see Fig. 4), we now take the limit $\epsilon\to 0$ with $\tilde{J}-1\sim\epsilon$ and $\tilde{K}\sim\epsilon$. At first order in $\epsilon$, Eqs. (V.2.2) and (34) become
$\partial_{t}r$ $=\left(\tilde{J}-1\right)r-\frac{1}{4}\left(1+\frac{1}{3}\cos(4\theta)\right)r^{3}\;,$ (40a)
@@
$\partial_{t}\theta=-\tilde{K}+\frac{\sin(4\theta)}{3+\cos(4\theta)}\left(\tilde{J}-1\right).$ (42)
Since $\sin(4\theta)/(3+\cos(4\theta))$ is bounded between $-1/(2\sqrt{2})$ and $1/(2\sqrt{2})$, a stationary state can only exist if ${\tilde{K}>(\tilde{J}-1)/(2\sqrt{2})}$. Therefore, in the regime $\tilde{J}-1\ll 1$, the SNIC bifurcation line corresponds to
$\tilde{K}_{\rm SNIC}=\frac{\tilde{J}_{\rm SNIC}-1}{2\sqrt{2}}.$ (43)
Note that, in Fig. 4, the thick blue line corresponding to the SNIC bifurcation was obtained numerically for the full nonlinear MF dynamics Eq. (24). Thus, our Eq. (43) only predicts the tangent of this thick blue line close to the bifurcation point $\tilde{J}_{\rm PF}=1$.
Finally, we find the stable points emerging at the SNIC bifurcation by extremizing $\sin(4\theta)/(3+\cos(4\theta))$ with respect to $\theta$. Those are
@@
for $n=0,1,2,3$, and their corresponding $r$ can be obtained from Eq. (41) as
@@
Approaching the SNIC bifurcation from the side of the limit cycle (swap phase), $\partial_{t}\theta$ approaches zero when the system passes at $\theta=\theta_{n}$. As a result, the oscillation’s period diverges at the bifurcation (see Fig. 5d-f, where the period increases as we approach the bifurcation). This stands in contrast with the Hopf bifurcation in which the period of oscillations remains finite.
@@ VI Monte-Carlo simulations
To go beyond the MF description, we perform large-scale Monte-Carlo simulations of the NR Ising model in both two and three dimensions (results in one dimension are summarized in Appendix D).
@@
Since three-dimensional simulations are much more costly, in 3D we update spins using a checkerboard algorithm to implement Eq. (46) instead of random picking. In the checkerboard algorithm non-interacting spins are flipped in parallel [152]. To generalize the checkerboard approach to the two-species case, we defined the quantity $\mathcal{P}=x_{1}+x_{2}+x_{3}+\alpha$ where $x_{n}$ is the integer position in the $n$-th dimension of the cubic lattice, and $\alpha=0,1$ labels the two species. Spins with the same parity of $\mathcal{P}$ (even or odd) are then updated simultaneously. This allows us both to parallelize the simulation and avoid random number generation in choosing the spins. We did not observe qualitative changes when changing the update rules from random picking to paralleled checkerboard, as well as when using other update rules, see Appendix A.
Unless mentioned otherwise, we initialize the system in an ordered state where all spins of the same species are either up or down, and let the system evolve until a steady-state is reached. Time is measured in simulation sweeps, where a single sweep is defined as $2\times L^{d}$ (the number of spins) MC steps.
@@
quantifies the time-dependence of the magnetizations $M_{A}$ and $M_{B}$. It is related to the entropy production [153] close to the Hopf bifurcation. In mean-field, the synchronization order parameter defined in Eq. (47) is
$R_{\rm MF}=\frac{r}{\sqrt{2}}$ (50)
@@
$\mathcal{L}_{\rm MF}=-r^{2}\dot{\theta}.$ (51)
The sign is chosen so that $\mathcal{L}_{\rm MF}$ is positive when $K$ is positive. In Fig. 6, $R_{\rm MF}$ and $\mathcal{L}_{\rm MF}$, calculated from a numerical stationary solution of Eq. (30), are plotted as a function of $\tilde{J}$ for a fixed value of $\tilde{K}$. It is shown that $R_{\rm MF}$ goes from zero to non-zero values when crossing the Hopf bifurcation at $\tilde{J}=1$. When crossing the SNIC bifurcation, $R_{\rm MF}$ remains finite, but there is a sharp change in the slope: the slope increases when approaching the bifurcation and decreases in the static order regime. Conversely, $\mathcal{L}_{\rm MF}$ is identically zero in both the disordered and static-order phases, but deviates from zero in the swap phase. We now analyze how these MF results hold in finite-dimensional systems using the Monte-Carlo simulations.
Fig. 6: Mean-field order parameters. (a) The synchronization order parameter $R_{\rm MF}$ (Eq. (50)) and (b) the phase-space angular momentum $\mathcal{L}_{\rm MF}$ (Eq. (51)), shown as a function of $\tilde{J}$ and for fixed $\tilde{K}=0.1$. The variables $r(t)$ and $\theta(t)$ are calculated from numerical solutions of Eq. (30) in steady state.

### v2

@@ V.1 Connection with locally-coupled oscillators
with the sum running over nearest-neighbor pairs. Equations (2729) can be viewed as describing locally-coupled oscillators with internal frequency and discrete states. In this sense, they share similarities with the Wood et al. cyclic model [95, 94, 96] although oscillators are restricted to forward moves (backward moves are prohibited) in these references. For $K=0$, our model reaches an equilibrium state that is equivalent to the clock model (vector Potts model) for $q=4$ [167, 168].
@@ V.2 Analysis of the bifurcations
In the ${m_{\alpha}\ll 1}$ limit and for small gradients, the mean-field equation in the fully anti-symmetric case (Eq. (25)) can be approximated by
@@
For $\tilde{K}=0$, Eq. (30) reduces to two uncoupled mean-field equations that derive from a $\phi^{4}$ Ginzburg-Landau potential. For any $\tilde{K}\neq 0$, Eq. (30) does not derive from a potential, and describes a non-equilibrium system. Note that due to the lack of rotational symmetry in the cubic term, Eq. (30) differs from the Complex Ginzburg-Landau (CGL) equation [86, 169], which is given by
@@
where $\psi({\vec{r}},t)$ is a complex field and $a$, $b$ and $c$ are complex numbers. However, Eq. (30) becomes CGL-like in limits discussed below.
Here, we restrict our analysis to spatially homogeneous solutions, $m_{\alpha}({\vec{r}},t)\equiv m_{\alpha}(t)$. The linear stability of such solutions to Eq. (30) is unaffected by the Laplacian term, and we therefore analyze the spatially homogeneous mean-field equation given by
@@
Figure 3 shows the phase-space velocity field $\vec{v}=(\partial_{t}m_{A},\partial_{t}m_{B})$ from the spatially homogeneous full mean-field equation, Eq. (25) along with the fixed points which are obtained by numerically solving the $\partial_{t}m_{A}=\partial_{t}m_{B}=0$ equations. The stability of the fixed points (shown in the figure as orange, pink, and brown for stable, saddle, and unstable, respectively) is determined by analyzing the eigenvalues of the Jacobian matrix $\mathcal{J}$ of the linearized equation $\partial_{t}\delta\vec{m}=\mathcal{J}\delta\vec{m}$ where linearization is performed around each fixed point.
@@
The mean-field phase diagram obtained from the spatially homogeneous full mean-field equation is shown in Fig. 4, as a function of the couplings $\tilde{J}$ and $\tilde{K}$. The phases are determined by the fixed points and their stability as described above. When nonreciprocal interactions are turned off (${\tilde{K}=0}$), a pitchfork bifurcation at ${\tilde{J}=1}$ (red point in Fig. 4) separates the disordered phase (in blue) from the static-order phase (in green), as in the usual Ising model. When nonreciprocal interactions are present (${\tilde{K}\neq 0}$), the swap phase arises (in pink) and is separated from the disordered state by a Hopf bifurcation at $\tilde{J}=1$ (green line), and from the static-order phase by another bifurcation known as a ($C_{4}$-symmetric) saddle-node on an invariant circle (SNIC) bifurcation [83, 64] (blue line). We analyze each of these bifurcations below using Eq. (32).
Fig. 4: Mean-field phase diagram. The three phases shown on the diagram are (i) disorder, (ii) swap and (iii) static order. The lines separating the phases are Hopf bifurcation (thin green) and SNIC bifurcation (thick blue). Red dot is the Pitchfork bifurcation. Obtained from numerical analysis of the spatially homogeneous Eq. (25).
@@ V.2.1 Pitchfork bifurcation
For ${\tilde{K}=0}$, Eq. (32) leads to the standard pitchfork bifurcation at ${\tilde{J}=1}$. The ${m_{A}=m_{B}=0}$ point becomes unstable and four stable fixed points
$m_{A}=\pm m_{B}=\pm\sqrt{\frac{3(\tilde{J}-1)}{\tilde{J}^{3}}}$ (33)
@@ V.2.2 Hopf bifurcation
This means that there is a Hopf bifurcation at ${\tilde{J}=1}$: the ${m_{A}=m_{B}=0}$ point become unstable for ${\tilde{J}}>1$ and a stable limit cycle emerges, corresponding to the swap phase.
The limit cycle is more easily analyzed in polar coordinates, ${m_{A}=r\cos\theta}$ and ${m_{B}=r\sin\theta}$ where Eq. (32) becomes
$\partial_{t}r$ $=(\tilde{J}-1)\,r-\frac{1}{4}\tilde{J}^{3}\bigg{[}1+\kappa^{2}$
@@
$-\left(\kappa-\frac{\kappa^{3}}{3}\right)\cos(4\theta)+\left(\frac{1}{3}-\kappa^{2}\right)\sin(4\theta)\bigg{]}r^{2}\;,$ (36)
@@
The limit cycle describes anharmonic oscillations, where both $r$ and $\dot{\theta}$ vary with $\theta$. Its shape has a rotational symmetry of order 4 due to the $C_{4}$ symmetry of the mean-field equation. However, very close to the bifurcation, i.e. $\tilde{J}-1\ll 1$ and any fixed $\tilde{K}$, the oscillations are harmonic [85]. This can be seen as follows.
@@
where we used the fact that ${r^{2}\sim\tilde{J}-1}$. The angle $\theta$ then has the solution ${\theta=\theta_{0}-\tilde{K}t}$, and the equation for $r$ becomes
$\partial_{t}r$ $=(\tilde{J}-1)\,r-\frac{1}{4}\bigg{[}1+\tilde{K}^{2}+\left(\frac{1}{3}-\tilde{K}^{2}\right)\cos(4\tilde{K}t)$ (38)
@@
$r\left(t\right)=2\sqrt{\frac{\tilde{J}-1}{1+\tilde{K}^{2}}}\times\frac{1}{\sqrt{1-(\tilde{J}-1)f(\tilde{J},\tilde{K},t)}}$ (39)
@@
$f(\tilde{J},\tilde{K},t)$ $=\frac{1-\tilde{J}+3(\tilde{J}-3)\tilde{K}^{2}+2\tilde{K}^{4}}{3(1+\tilde{K}^{2})\left[(\tilde{J}-1)^{2}+4\tilde{K}^{2}\right]}\cos(4\tilde{K}t)$ (40)
$-\frac{\tilde{K}\left(5-3\tilde{J}+(\tilde{J}-7)\tilde{K}^{2}\right)}{3(1+\tilde{K}^{2})\left[(\tilde{J}-1)^{2}+4\tilde{K}^{2}\right]}\sin(4\tilde{K}t)$
@@
which is a circular motion of constant angular velocity in phase space. This harmonic solution can be rationalized by looking at the two times scales corresponding respectively to the angular and radial motion near the bifurcation (described by Eqs. (37) and (38), respectively). The angular motion time scale is $T_{\theta}\sim\tilde{K}^{-1}$ while the radial motion time scale is determined by the relaxation time of $r(t)$ to its new minimum when $\theta$ changes, which scales as $T_{r}\sim(\tilde{J}-1)^{-1}$. The two time scales $T_{\theta}$ and $T_{r}$ can be interpreted as “oscillation time-scale” and “alignment time-scale”, respectively. Since $T_{r}\gg T_{\theta}$, the angular motion occurs much more rapidly, leaving insufficient time for $r$ to reach a new minimum when $\theta$ changes. Thus, only the $\theta$-averaged value of the right-hand side of Eq. (V.2.2) matters. This is supported by Fig. 5 which shows typical phase-space trajectories in the swap phase: as $\tilde{J}-1$ increases, the motion indeed loses its rotational symmetry and become more “square-like”.
Fig. 5: Mean-field swap phase at different distances from the Hopf and SNIC bifurcations. Numerical solutions of the spatially homogeneous mean-field dynamics, Eq. (32). (a-c) $m_{B}$ vs. $m_{A}$. (d-f) $m_{A}$ vs. $t$. From left to right: $\tilde{J}=1.0002$, $1.001$, and $1.0028$. In all plots $\tilde{K}=0.001$. Red dot is the initial state and dashed black line is the steady state predicted from Eq. (41), valid near the Hopf bifurcation.
The above result is a consequence of a more general theorem. A reaction-diffusion equation that undergoes a Hopf bifurcation such as the mean-field evolution Eq. (30) can be reduced, close to the bifurcation, to the CGL equation (i.e., the CGL equation is its normal form). The transformation to the CGL equation guarantees that the steady-state solution of the non-transformed set of equations is an elliptical limit cycle (or in special cases like here, circular) [85]. Because of this reduction, we expect the disorder-to-swap phase transition in our model to be in the same universality class as the Hopf bifurcation in the CGL equation, which has been studied in Refs. [89, 90].
@@ V.2.3 Infinite-period (SNIC) bifurcation
Let us now start from the limit-cycle swap phase. Upon increasing $\tilde{J}$ or decreasing $\tilde{K}$, the system goes through a ($C_{4}$-symmetric) saddle-node on an invariant circle (SNIC) bifurcation, where four pairs of stable and saddle fixed points emerge from the limit cycle (pink and orange points in Fig. 3c). The SNIC bifurcation is also known under several other names: infinite-period bifurcation (because the period of oscillation diverges at the bifurcation point), saddle-node on a limit cycle (SNLC), saddle-node infinite-period (SNIPER), etc. As the dynamics is invariant under the action of the group $C_{4}$ (see Sec. V), four copies of the SNIC bifurcation happen at the same time. Bifurcations in such a dynamical system with symmetries are known as equivariant bifurcations, see Refs. [170, 171, 172] for an introduction. Once the stable fixed points emerge, the system is in a static-order phase.
Noticing that the SNIC bifurcation occurs when $\tilde{K}$ and $\tilde{J}-1$ are comparable (see Fig. 4), we now take the limit $\epsilon\to 0$ with $\tilde{J}-1\sim\epsilon$ and $\tilde{K}\sim\epsilon$. At first order in $\epsilon$, Eqs. (V.2.236) become
$\partial_{t}r$ $=(\tilde{J}-1)\,r-\frac{1}{4}\left(1+\frac{1}{3}\cos(4\theta)\right)r^{3}\;,$ (42a)
@@
$\partial_{t}\theta=-\tilde{K}+\frac{\sin(4\theta)}{3+\cos(4\theta)}(\tilde{J}-1).$ (44)
Since $\sin(4\theta)/(3+\cos(4\theta))$ is bounded between $-1/(2\sqrt{2})$ and $1/(2\sqrt{2})$, a stationary state (with $\partial_{t}\theta=0$) can only exist if ${\tilde{K}<(\tilde{J}-1)/(2\sqrt{2})}$. Therefore, in the $\tilde{J}-1\ll 1$ regime, the SNIC bifurcation line corresponds to
$\tilde{K}=\frac{\tilde{J}-1}{2\sqrt{2}}.$ (45)
Note that, in Fig. 4, the thick blue line corresponding to the SNIC bifurcation was obtained numerically for the full mean-field dynamics Eq. (25). Equation (45) only predicts the tangent of this thick blue line close to the pitchfork bifurcation point ($\tilde{J}=1$, $\tilde{K}=0$).
Finally, the fixed points emerging at the SNIC bifurcation are obtained from the $\theta$ values that maximize $\sin(4\theta)/(3+\cos(4\theta))$, which are
@@
with $n=0,1,2,3$. Together with their corresponding $r$ values derived from Eq. (43) we obtain
@@
Approaching the SNIC bifurcation from the side of the limit cycle (swap phase), $\partial_{t}\theta$ approaches zero when the system passes at $\theta=\theta_{n}$. As a result, the oscillation period diverges at the bifurcation (see Fig. 5d-f, where the period increases as we approach the bifurcation). This stands in contrast with the Hopf bifurcation in which the oscillation period remains finite.
@@ VI Monte-Carlo simulations
To go beyond the mean-field description, we perform large-scale Monte-Carlo simulations of the nonreciprocal Ising model in both two and three dimensions (results in one dimension are summarized in Appendix B).
@@
Since three-dimensional simulations are much more costly, in 3D we update spins using a checkerboard algorithm to implement Eq. (48) instead of random picking. In the checkerboard algorithm non-interacting spins are flipped in parallel [173]. To generalize the checkerboard approach to the two-species case, we defined the quantity $\mathcal{P}=x_{1}+x_{2}+x_{3}+\alpha$ where $x_{n}$ is the integer position in the $n$-th dimension of the cubic lattice, and $\alpha=0,1$ labels the two species. Spins with the same parity of $\mathcal{P}$ (even or odd) are then updated simultaneously. This allows us both to parallelize the simulation and avoid random number generation in choosing the spins. We did not observe qualitative changes when changing the update rules from random picking to paralleled checkerboard, as well as when using other update rules, see Appendix C.
Unless mentioned otherwise, we initialize the system in an ordered state where all spins of the same species are either up or down, and let the system evolve until a steady state is reached. Averages are then calculated in the steady state, excluding the initial transient. Time is measured in simulation sweeps, where a single sweep is defined as $2\times L^{d}$ (the number of spins) MC steps (individual attempted flips). To best visualize phenomena of interest, simulation snapshots are presented either with the spin variables $\sigma^{A}$ and $\sigma^{B}$, or with the angle variable $\theta$.
@@
quantifies the macroscopic time-reversal symmetry breaking. It is related to the entropy production [174], see Appendix A for details.
Mean-field behavior of the order parameters
In mean-field and for a spatially uniform state, the synchronization order parameter defined in Eq. (49) is
$R_{\rm MF}=\frac{\langle r\rangle_{t}}{\sqrt{2}}$ (52)
@@
$\mathcal{L}_{\rm MF}=-\langle r^{2}\dot{\theta}\rangle_{t}.$ (53)
The sign is chosen so that $\mathcal{L}$ is positive when $K$ is positive. In Fig. 6, $R_{\rm MF}$ and $\mathcal{L}_{\rm MF}$, calculated from a numerical steady-state solution of the spatially homogeneous Eq. (25), are plotted as a function of $\tilde{J}$ for a fixed value of $\tilde{K}$. It is shown that $R_{\rm MF}$ goes from zero to non-zero values when crossing the Hopf bifurcation at $\tilde{J}=1$. When crossing the SNIC bifurcation, $R_{\rm MF}$ remains finite, but there is a sharp change in the slope: the slope increases when approaching the bifurcation and decreases in the static order regime. Conversely, $\mathcal{L}_{\rm MF}$ is identically zero in both the disordered and static-order phases, but deviates from zero in the swap phase. Both $R_{\rm MF}$ and $\mathcal{L}_{\rm MF}$ are continuous at the bifurcation points, as in a second-order phase transition. In the swap phase and near the Hopf bifurcation, we can use Eq. (37) and Eq. (41) to obtain
$R_{\rm MF}=\sqrt{\frac{2(\tilde{J}-1)}{1+\tilde{K}^{2}}}\quad\text{and}\quad\mathcal{L}_{\rm MF}=\frac{4\tilde{K}(\tilde{J}-1)}{1+\tilde{K}^{2}}.$ (54)
In the swap phase and near the SNIC bifurcation, $\mathcal{L}_{\rm MF}$ scales with $\mu^{1/2}$ where $\mu$ is the distance from the bifurcation line (because the period scales as $T\sim\mu^{-1/2}$ while $r$ remains finite at the bifurcation [83]). We now analyze how these mean-field results hold in finite-dimensional systems using the Monte-Carlo simulations.
Fig. 6: Mean-field order parameters. (a) The synchronization order parameter $R_{\rm MF}$ (Eq. (52)) and (b) the phase-space angular momentum $\mathcal{L}_{\rm MF}$ (Eq. (53)), shown as a function of $\tilde{J}$ and for fixed $\tilde{K}=0.1$. Black lines are obtained from numerical solutions of the spatially homogeneous Eq. (25) in steady state. Dashed green lines are approximate solutions that are valid near the Hopf bifurcation, given by Eq. (54).